An electron wave making a standing wave in a hydrogen atom has a wavelength of 8.94 × 10−11 m. If the mass of an electron is 9.11 × 10−31 kg, what is the velocity of the electron according to de Broglie equation?
Answer #1

given

= wavelength of 8.94 x 10−11 m

m = mass of an electron is 9.11 x 10−31 kg

the velocity of the electron according to de Broglie equation is v = ?

using de Broglie equation = h / m v

here h = planck's constant

= 6.626 x 10-34 J.sec

= h / m v

8.94 x 10−11 = 6.626 x 10-34 / 9.11 x 10−31 x v

v = 6.626 x 10-34 / 9.11 x 10−31 x 8.94 x 10−11

v = 8135711.427 m/sec

or

the velocity of the electron according to de Broglie equation is v = 8.135711427 x 106 m/sec